OpenStudy (anonymous):

Subtract and give the answer in simplest form:

OpenStudy (anonymous):

\[3x \over x-5\] - \[5 \over 5-x\]

OpenStudy (anonymous):

Answer choices: A. 3 B. \[3x-5 \over x-5\] C. \[3x+5 \over x-5 \]

OpenStudy (anonymous):

3x(5-x)- 5(x-5) (3x+5)(5-x) = -------------- = ---------- (x-5)(5-x) (x-5)(5-x) = 3x+5 ---- x - 5

OpenStudy (anonymous):

How did you group (5-x) and (x-5)???

OpenStudy (anonymous):

(5-x)(x-5) is the multiple of 5-x and x-5

OpenStudy (anonymous):

you multiply both original fractions by (5-x)(x-5)

OpenStudy (anonymous):

I know, but you can't group it because it's different we can group (5-x) and (5-x), not (5-x)(x-5) your answer is wrong

OpenStudy (anonymous):

ok - -i'll check it out with pen and paper

OpenStudy (anonymous):

its correct

OpenStudy (anonymous):

oh, I'm stupid \[3x(5-x)-5(x-5)=3x(5-x)-5(-1)(5-x)=3x(5-x)+5(5-x)=(5-x)(3x+5)\] yes, it's correct)

OpenStudy (anonymous):

lol - human error - i was like this a few days ago - only worse